# Gaussian Copula-Based Outage Performance Analysis of Fluid Antenna Systems: Channel Coefficient- or Envelope-Level Correlation Matrix?

### v1

@@ V Conclusions
This paper explored the modeling accuracy of using envelope-level correlation matrix under Gaussian copula framework, and developed a method to generate fully correlated Nakagami-$m$ fading channel envelopes, enabling exact Monte Carlo simulations for validating the theoretical results. Simulation results confirm that using the envelope-level correlation matrix yields higher modeling accuracy, especially in sparse port deployment and low-outage regime. Moreover, as the average transmit SNR increases, Gaussian copula using the envelope-level correlation matrix yields more accurate OP than using coefficient-level one under denser port deployment. As the Nakagami-$m$ fading parameters $m$ decreases or $\mu$ increases, envelope correlation-based Gaussian copula achieves higher accuracy than coefficient at lower average transmit SNR. Building on these, in ultra-reliable communication or sparse port deployment scenarios, Gaussian copula with envelope-level correlation matrix provides an effective approach. Moreover, this approach can be further validated under other fading channel models, such as cascaded channels.

### v3

@@ V Conclusions
This paper explored the modeling accuracy of using envelope-level correlation matrix under Gaussian copula framework, and developed a method to generate fully correlated Nakagami-$m$ fading channel envelopes, enabling exact Monte-Carlo simulations for validating the theoretical results. Simulation results confirm that using the envelope-level correlation matrix yields higher modeling accuracy, especially in sparse port deployment and low-outage regime. Moreover, as the average transmit SNR increases, Gaussian copula using the envelope-level correlation matrix yields more accurate OP than using coefficient-level one under denser port deployment. As the Nakagami-$m$ fading parameters $m$ decreases or $\mu$ increases, envelope correlation-based Gaussian copula achieves higher accuracy than the coefficient correlation matrix. Building on these, in ultra-reliable communication or sparse port deployment scenarios, Gaussian copula using the envelope-level correlation matrix provides an effective performance evaluation approach. Moreover, it can be further validated under other fading channel models, such as cascaded channels.